OpenStudy (anonymous):

Which of the following is an example of the associative property of multiplication? 7 • (3 + 5) = 21 + 35 7 • (3 • 5) = (7 • 3) • 5 7 • (3 • 5) = 7 • (5 • 3) 7 • = 1

OpenStudy (anonymous):

The general example is given as "a + (b + c) = (a + b) + c"; a numeral example might be "7 + (3 + 14) = (7 + 3) + 14".